MICROWAVE SPECTRA OF THE O_{2}-HF COMPLEX

\underline{ERIK M. GRUMSTRUP}, GALEN SEDO and KENNETH R. LEOPOLD, Department of Chemistry, University of Minnesota, Minneapolis, MN 55455.

Microwave spectra have been recorded for the complex O_{2}-HF. Spectra were readily located based on results of previous infrared work.^{a} but yield spectroscopic constants of somewhat higher accuracy. The observed transitions show well resolved structure arising from the ^{1}H and ^{1}^{9}F nuclear spins. Magnetic super-hyperfine structure due to the interaction of the proton and fluorine nuclei with the spin magnetic moment of O_{2} appear to be of comparable magnitude to the HF spin-spin interaction in at least some of the observed transitions. Progress on the simultaneous analysis of these hyperfine and super-hyperfine effects will be reported.

aW. A. Fawzy, C. M. Lovejoy, D. J. Nesbitt, and J. T. Hougen J. Chem. Phys. \underline{117}(2), 693 (2002).